# Token Distillation: Attention-Aware Input Embeddings for New Tokens

### v2

@@ 1 Introduction
We compare against strong baselines, including standard embedding initialization procedures, training of the embedding matrices with causal language modeling, as well as pretrained hyper-networks that predict new token embeddings. Extensive experiments on a wide range of open-weight models, including question-answering benchmarks and the generation of definitions for the newly added tokens confirm the effectiveness of our approach. Our experimental setup is detailed in Section 4. Additionally, we describe related work in Section 2 and explicitly discuss limitations of our method in Limitations. In summary, our contributions are as follows.
@@ 3.3 Further Details
Like other methods optimizing embeddings based on contexts, our method also needs input sequences. Firstly, we note that randomly sampling texts from a domain-specific or general corpus is inefficient for our goal of learning newly added embeddings: typically, the new tokens will only make up a small fraction, so most gradient updates will actually not affect the new input embeddings at all and simply learn to minimize their log-probability in the case of output embeddings. As we aim to have a fast method, we need a better approach. We propose two different approaches: (1) Our main approach is to simply retrieve snippets that contain our target tokens from a domain-specific or general corpus. This can be implemented efficiently using the algorithm proposed by Aho & Corasick (1975). Then we can truncate the snippets to a small window around our target token to optimize computational efficiency. (2) For causal language models, we can simply generate snippets by prompting the model with our target token. We provide implementation details in Section B.7. In our main experiments, we focus on the first approach but note the availability of the second approach in case a reference corpus is not available. In most cases of domain and language adaptation, the availability of such a corpus is a reasonable assumption. Nevertheless, we study an ablation instead using the generative approach in Section 5.2.
@@
We employ a simple setup and use the AdamW (Kingma & Ba, 2017; Loshchilov & Hutter, 2019) optimizer for all trainable parameters. We set the batch size to optimize throughput and run all experiments on Nvidia H100 80GB GPUs. We do not use weight decay and maintain a constant learning rate with a linear warmup. For fair comparison, we sweep for the best learning rate for all methods that require a learning rate. Since our method is aimed to serve as an “initialization” rather than as full-scale further training, we restrict the number of example sequences to a maximum of 25 per target token and truncate to a context length of 50 tokens. These restrictions ensure that our method is quick to run, initializing 2,500 new tokens on a single GPU in under 10 minutes. We use the same data for all training-based methods, including baselines and variations of our method. We provide all details in Appendix B.
@@ 4.1 Evaluation
Furthermore, to push the limits of embedding initialization methods, we move beyond domain and language adaptation and apply our method to multi-word tokens, which have recently gained further interest (Gee et al., 2023; Liu et al., 2025; Huang et al., 2025). We follow the zero-shot tokenizer transfer setting in our evaluations (Minixhofer et al., 2024), i.e., we evaluate initialization methods without further training on a corpus, as we want to directly judge the quality of the resulting representations.
We select target tokens that actually occur frequently in the chosen benchmarks so that we can effectively judge the quality of newly generated embeddings. We select words that occur more frequently than a threshold and additionally ensure that all individual benchmarks are represented. We report the full methodology in Section B.1. For our multi-word token experiments, we prompt gpt-o4-mini-high to generate suitable candidates (full details in Section B.3). For our experiments in the biomedical domain, we retrieve contexts from ncbi/pubmed as a reference corpus. For our experiments on language adaptation to French, we use HuggingFaceFW/fineweb-2. For our multi-word token experiments, we prompt the original models to generate sequences containing the new tokens. For the LLM-as-a-Judge evaluations, we use Llama3.3-70B-Instruct as the judge model. We evaluate the general correctness of the generated definitions to test the quality and completeness of the resulting representations as well as semantic similarity with the target model’s output (Villegas et al., 2025), as inducing as little behavior change as possible can be an important desideratum.
@@ 4.2 Modeling
We conduct experiments using a wide range of open-weight model checkpoints to ensure our results are not merely a function of peculiarities in any specific model’s embedding space. In particular, we choose the following models: Mistral-7B-v0.1, Llama3-8B, Llama3-8B-Instruct, Llama3.1-8B, Llama3.1-8B-Instruct, Llama3.2-3B, Llama3.2-3B-Instruct, and OLMo2-7B-1124-Instruct. In the remainder of the paper, we denote “-Instruct” variants with a simple “-i” suffix. Through this extensive evaluation, we study different model families, model sizes (3B, 7B, and 8B), instruction-tuned and base models, as well as models with separate input/output embeddings and tied (shared) embeddings between input and output. For our experiments on French language adaptation, we run experiments with the Mistral-7B-v0.1 and Llama3-8B(-Instruct) models and additionally consider Qwen3-8B-Base, which has received more extensive multilingual training.
@@
Since our proposed method conducts a short optimization on a few reference sequences per new token, we compare against the common approach of training embeddings using causal language modeling using the same data. Specifically, we report results for “classic” embedding tuning using the next-token prediction objective – denoted as NTP (tune all embeddings) – as well as masking updates to the original embeddings such that only new embeddings are optimized (denoted as just NTP), which corresponds to the method used by Lampinen & McClelland (2018). These methods use the subtoken mean as their starting point. A strong alternative method for obtaining embeddings for new tokens are hyper-networks specifically pretrained for this task. We report results for ZeTT (Minixhofer et al., 2024) using their provided hyper-network checkpoints. We also compare against “Tokens to Words” (Kaplan et al., 2025), which relies on PatchScopes (Ghandeharioun et al., 2024) to extract hidden states and learns mappings from these to the embedding spaces. Unlike our Token Distillation approach, it hinges on a successful PatchScopes extraction of hidden states for each particular token. In our experiments, “Tokens to Words” could not be applied to 10–75% of tokens due failed PatchScope extractions, skewing the comparison with other methods that do not suffer from this limitation. We therefore report its results separately (Section A.1); Token Distillation performs better on average while also covering all added tokens. We run five random seeds for all methods that include randomness and report the mean and standard deviation.
@@
As the initial embedding input into our method, we also use the subtoken mean. We investigate alternatives to this choice in Section 5.2. Our proposed objective can also be combined with the NTP-based objectives. We report results for combining Token Distillation and NTP. Since the NTP and Token Distillation objectives have different scales, a simple addition is suboptimal. In fact, NTP is usually of larger magnitude, which leads it to overpower Token Distillation. Therefore, we also consider an “autoscaled” variant Token Distillation + NTP where the NTP loss is scaled by = stop_gradient(loss_awedist / loss_ntp) before summing. Here, stop_gradient prevents gradient flow through the scaling factor .333Multi-objective optimization is a rich field with extensive literature, we leave a thorough investigation of alternatives such as GradNorm (Chen et al., 2018) for future work.
@@
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 Random (Hewitt, 2021) 57.0±0.6 58.8±0.5 60.6±0.3 59.3±0.5 62.6±0.6 51.6±0.4 55.8±0.2 53.9±0.6 57.5 NTP (tune all embeddings) 55.7±0.5 63.8±0.1 63.8±0.2 62.9±0.4 66.8±0.4 51.5±0.6 55.5±0.5 58.0±0.3 59.8 Mean (Gee et al., 2022) 58.3 63.8 63.4 63.7 66.9 54.2 58.4 58.0 60.8 NTP (Lampinen et al., 2018) 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 ZeTT (Minixhofer et al., 2024) 62.7 66.1 66.3 – – – – – – Token Distillation 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 Token Distillation + NTP 63.0±0.5 67.2±0.3 66.7±1.6 67.2±0.3 70.6±0.4 57.6±0.3 62.2±0.3 59.8±0.5 64.3 Token Distillation + NTP 62.8±0.5 67.6±0.4 67.8±0.5 67.4±0.4 70.9±0.3 57.9±0.1 62.5±0.6 61.2±0.2 64.7
@@
Method Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Original tokenization 99.8 96.5 100.0 94.3 100.0 98.4 100.0 94.7 100.0 98.4 100.0 93.2 100.0 93.8 99.4 96.9 99.9 95.8 Random (Hewitt, 2021) 0.0 0.4 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.1 Mean (Gee et al., 2022) 2.5 4.3 16.8 16.2 17.6 20.7 25.6 25.6 23.2 28.3 15.2 15.2 12.1 14.6 20.1 24.2 16.6 18.6 NTP (tune all embeddings) 26.8 30.3 33.2 39.5 34.4 43.4 28.5 35.0 41.4 50.0 19.7 20.7 18.6 21.1 22.5 26.6 28.1 33.3 NTP (Lampinen et al., 2018) 49.8 58.2 45.3 49.6 48.4 58.0 58.6 65.6 60.2 69.9 51.6 58.4 50.0 58.8 52.1 57.0 52.0 59.4 ZeTT (Minixhofer et al., 2024) 63.5 68.6 69.5 73.6 70.7 80.3 – – – – – – – – – – – – Token Distillation 79.7 85.2 72.7 76.2 79.7 91.0 75.8 80.1 81.6 89.8 0.0 0.0 75.0 83.0 83.8 89.5 68.5 74.4 Token Distillation + NTP 77.3 82.6 68.4 73.8 75.8 86.3 72.3 79.9 77.3 86.1 62.3 69.1 65.2 75.8 65.4 74.6 70.5 78.5 Token Distillation + NTP 76.8 82.0 71.7 75.2 80.9 91.6 77.5 81.2 81.1 89.5 72.7 79.5 71.9 80.5 81.2 86.9 76.7 83.3
@@
Mistral-7B Llama3-8B Llama3-8B-i Qwen3-8B Avg. Original tokenization 69.5 69.4 72.1 81.7 73.2 Random (Hewitt, 2021) 51.0 46.1 51.5 62.1 52.7 Mean (Gee et al., 2022) 56.3 58.4 61.7 69.6 61.5 NTP (Lampinen et al., 2018) 64.7 67.0 70.1 81.2 70.8 ZeTT (Minixhofer et al., 2024) 66.1 61.3† 65.2† – – Token Distillation 68.5 68.9 72.9 81.5 72.9
@@ 5.1 Main Experiments
Note that even with its very competitive performance, Token Distillation does not fully attain the level of the original tokenization, which benefits from the massive pretraining of the original model. This is expected in our zero-shot tokenizer transfer setting without any further training (see Minixhofer et al., 2024), although remarkably, Token Distillation is able to outperform the original tokenization when adding French tokens to Llama3-8B-i without any training of the Transformer layers. We seek to directly investigate the embedding initialization quality and the advantages of our distillation-based objective compared to next-token prediction rather than the effects of further training. However, results can naturally be improved via further training; we argue that Token Distillation will provide the best starting point for such further processing. Additionally, we show in Appendix C that an additional epoch of further Token Distillation is able to improve results, whereas an additional epoch with the next-token prediction objective plateaus. In the following, we analyze various aspects of the results in further detail.
Note that in Table 2, only for Llama3.2-3B, Token Distillation has performance on par with random initialization. Llama3.2-3B(-i) are the only models in our lineup with tied embedding weights. Our objective does not explicitly enforce a bound on the norm of the new embedding, which in this case led to the failure mode of always generating a specific new embedding with very large norm. Note that this does not always happen for checkpoints with tied weights, as evidenced by the results for Llama3.2-3B-i. However, we additionally propose a simple modification to alleviate the occurrence of this issue by combining Token Distillation with the next-token prediction objective, which (implicitly) acts as a regularizer on the output embedding norm. In support of our argument that a subtoken attention distillation objective is superior to next-token prediction for learning new token embeddings, the combination via a sum of the two objectives (Token Distillation + NTP) generally yields worse results than Token Distillation alone. However, we can add a dynamic downweighting factor (see Section 4.2) to the next-token prediction objective (Token Distillation + NTP), which mostly alleviates the negative interference while keeping the regularizing effect.
@@
Vanilla NTP-tuning of all embeddings (denoted as tune all embeddings) yields disappointing results, even underperforming the subtoken mean initialization. We investigate this and observe a degradation of the original token embeddings. Note that for Token Distillation, we only optimize new token embeddings. Therefore, we also compare against the NTP baseline, where we similarly optimize only the new token embeddings. NTP outperforms NTP, supporting our analysis. However, even with this improvement, Token Distillation in turn still further surpasses NTP.
@@
ZeTT, which uses a pretrained hypernetwork, is the strongest baseline we compare against. We only compare against ZeTT when a pretrained hyper-network by Minixhofer et al. (2024) is available444Minixhofer et al. (2024) do not provide a checkpoint trained specifically for the -Instruct variant of Llama3-8B but show that a transfer from the base version is possible.. We first note that ZeTT yields quite impressive results, outperforming even our optimized NTP baseline. Evidently, the pretraining of ZeTT’s hypernetwork has internalized a better embedding prediction than the iterative gradient descent optimization on selected samples we employ for NTP. Nevertheless, our proposed method Token Distillation outperforms even ZeTT – without needing any hyper-network pretraining. It is important to note that at inference time, hyper-network based methods such as ZeTT are actually faster than Token Distillation (and NTP-based baselines), since only a single forward pass through the hyper-network per token is required. However, they require expensive pretraining for each new model and – in some cases – target domain. For example, ZeTT with a hypernetwork pretrained only on English naturally underperforms on language adaptation to other languages – as this is out of distribution from the hypernetwork training, whereas the multilingually pretrained variant available for one of the models performs more favorably (see Table 3). We note that in concurrent work, Minixhofer et al. (2025) have proposed to train embedding prediction hyper-networks with a distillation-based objective and show improvements over the next-token prediction objective, which is in line with our findings. However, this does not yet solve the domain-specificity problem of pretrained hyper-networks. Alternatively, we can actually use ZeTT-generated embeddings as a starting point and further tune them using our method for even better results. We analyze this possibility in Section 5.2.

### v3

@@ 1 Introduction
We compare against strong baselines, including standard embedding initialization procedures, training of the embedding matrices with causal language modeling, as well as pretrained hyper-networks that predict new token embeddings. Extensive experiments on a wide range of open-weight models, including question-answering benchmarks and the generation of definitions for the newly added tokens confirm the effectiveness of our approach. Our experimental setup is detailed in Section 4. Additionally, we describe related work in Section 2 and explicitly discuss limitations of our method in Appendix A. In summary, our contributions are as follows.
@@ 3.3 Further Details
Like other methods optimizing embeddings based on contexts, our method also needs input sequences. Firstly, we note that randomly sampling texts from a domain-specific or general corpus is inefficient for our goal of learning newly added embeddings: Typically, the new tokens will only make up a small fraction, so most gradient updates will actually not affect the new input embeddings at all and simply learn to minimize their log-probability in the case of output embeddings. As we aim to have a fast method, we need a better approach. We propose two different approaches: (1) Our main approach is to simply retrieve snippets that contain our target tokens from a domain-specific or general corpus. This can be implemented efficiently using the algorithm proposed by Aho & Corasick (1975). Then we can truncate the snippets to a small window around our target token to optimize computational efficiency. (2) For causal language models, we can simply generate snippets by prompting the model with our target token. We provide implementation details in Section C.8. In our main experiments, we focus on the first approach but note the availability of the second approach in case a reference corpus is not available. In most cases of domain and language adaptation, the availability of such a corpus is a reasonable assumption. Nevertheless, we study an ablation instead using the generative approach in Section 5.3.
@@
We employ a simple setup and use the AdamW (Kingma & Ba, 2017; Loshchilov & Hutter, 2019) optimizer for all trainable parameters. We set the batch size to optimize throughput and run all experiments on Nvidia H100 80GB GPUs. We do not use weight decay and maintain a constant learning rate with a linear warmup. For fair comparison, we sweep for the best learning rate for all methods that require a learning rate. Since our method is aimed to serve as an initialization rather than as full-scale further training, we restrict the number of example sequences to a maximum of 25 per target token and truncate to a context length of 50 tokens. These restrictions ensure that our method is quick to run, initializing 2,500 new tokens on a single GPU in under 10 minutes. In Section 5.2, we additionally report experiments with a larger compute budget (100 snippets per token) and with continued training for vocabulary adaptation. We use the same data for all training-based methods, including baselines and variations of our method. We provide all details in Appendix C.
@@ 4.1 Evaluation
Furthermore, to push the limits of embedding initialization methods, we move beyond domain and language adaptation and apply our method to multi-word tokens, which have recently gained further interest (Gee et al., 2023; Liu et al., 2025; Huang et al., 2025). We follow the zero-shot tokenizer transfer setting in our evaluations (Minixhofer et al., 2024), i.e., we evaluate initialization methods without further training on a corpus, as we want to directly judge the quality of the resulting representations. In Section 5.2, we additionally experiment with a short continued training phase after initializing new embeddings.
We select target tokens that actually occur frequently in the chosen benchmarks so that we can effectively judge the quality of newly generated embeddings. We select words that occur more frequently than a threshold and additionally ensure that all individual benchmarks are represented. We report the full methodology in Section C.1. For our multi-word token experiments, we prompt gpt-o4-mini-high to generate suitable candidates (full details in Section C.3). For our experiments in the biomedical domain, we retrieve contexts from ncbi/pubmed as a reference corpus. For our experiments on language adaptation to French, we use HuggingFaceFW/fineweb-2. For our multi-word token experiments, we prompt the original models to generate sequences containing the new tokens. For the LLM-as-a-Judge evaluations, we use Llama3.3-70B-Instruct as the judge model. We evaluate the general correctness of the generated definitions to test the quality and completeness of the resulting representations as well as semantic similarity with the target model’s output (Villegas et al., 2025), as inducing as little behavior change as possible can be an important desideratum.
@@ 4.2 Modeling
We conduct experiments using a wide range of open-weight model checkpoints to ensure our results are not merely a function of peculiarities in any specific model’s embedding space. In particular, we choose the following models: Mistral-7B-v0.1, OLMo2-7B-1124-Instruct, Llama3-8B, Llama3-8B-Instruct, Llama3.1-8B, Llama3.1-8B-Instruct, Llama3.2-3B, and Llama3.2-3B-Instruct. In the remainder of the paper, we denote “-Instruct” variants with a simple “-i” suffix. Through this extensive evaluation, we study different model families, model sizes (3B, 7B, and 8B), instruction-tuned and base models, as well as models with separate input/output embeddings and tied (shared) embeddings between input and output. For our experiments on French language adaptation, we run experiments with the Mistral-7B-v0.1 and Llama3-8B(-Instruct) models and additionally consider Qwen3-8B-Base, which has received more extensive multilingual training.
@@
Since our proposed method conducts a short optimization on a few reference sequences per new token, we compare against the common approach of training embeddings using causal language modeling using the same data. Specifically, we report results for “classic” embedding tuning using the next-token prediction objective – denoted as NTP (tune all embeddings) – as well as masking updates to the original embeddings such that only new embeddings are optimized (denoted as just NTP), which corresponds to the method used by Lampinen & McClelland (2018). These methods use the subtoken mean as their starting point. A strong alternative method for obtaining embeddings for new tokens are hyper-networks specifically pretrained for this task. We report results for ZeTT (Minixhofer et al., 2024) using their provided hyper-network checkpoints.
@@
As the initial embedding input into our method, we also use the subtoken mean. We investigate alternatives to this choice in Section 5.3. Our proposed objective can also be combined with the NTP-based objectives. We report results for combining Token Distillation and NTP. Since the NTP and Token Distillation objectives have different scales, a simple addition is suboptimal. In fact, NTP is usually of larger magnitude, which leads it to overpower Token Distillation. Therefore, we also consider an “autoscaled” variant Token Distillation + NTP where the NTP loss is scaled by = stop_gradient(loss_token_distillation / loss_ntp) before summing. Here, stop_gradient prevents gradient flow through the scaling factor .333Multi-objective optimization is a rich field with extensive literature. We leave a thorough investigation of alternatives such as GradNorm (Chen et al., 2018) for future work.
@@
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 Random (Hewitt, 2021) 57.0±0.6 58.8±0.5 60.6±0.3 59.3±0.5 62.6±0.6 51.6±0.4 55.8±0.2 53.9±0.6 57.5 NTP (tune all embeddings) 55.7±0.5 63.8±0.1 63.8±0.2 62.9±0.4 66.8±0.4 51.5±0.6 55.5±0.5 58.0±0.3 59.8 Mean (Gee et al., 2022) 58.3 63.8 63.4 63.7 66.9 54.2 58.4 58.0 60.8 NTP Lampinen et al., 2018 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 ZeTT (Minixhofer et al., 2024) 62.7 66.1 66.3 – – – – – – Token Distillation 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 Token Distillation + NTP 63.0±0.5 67.2±0.3 66.7±1.6 67.2±0.3 70.6±0.4 57.6±0.3 62.2±0.3 59.8±0.5 64.3 Token Distillation + NTP 62.8±0.5 67.6±0.4 67.8±0.5 67.4±0.4 70.9±0.3 57.9±0.1 62.5±0.6 61.2±0.2 64.7
@@
Method Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Original tokenization 99.8 96.5 100.0 94.3 100.0 98.4 100.0 94.7 100.0 98.4 100.0 93.2 100.0 93.8 99.4 96.9 99.9 95.8 Random (Hewitt, 2021) 0.0 0.4 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.1 Mean (Gee et al., 2022) 2.5 4.3 16.8 16.2 17.6 20.7 25.6 25.6 23.2 28.3 15.2 15.2 12.1 14.6 20.1 24.2 16.6 18.6 NTP (tune all embeddings) 26.8 30.3 33.2 39.5 34.4 43.4 28.5 35.0 41.4 50.0 19.7 20.7 18.6 21.1 22.5 26.6 28.1 33.3 NTP Lampinen et al., 2018 49.8 58.2 45.3 49.6 48.4 58.0 58.6 65.6 60.2 69.9 51.6 58.4 50.0 58.8 52.1 57.0 52.0 59.4 ZeTT (Minixhofer et al., 2024) 63.5 68.6 69.5 73.6 70.7 80.3 – – – – – – – – – – – – Token Distillation 79.7 85.2 72.7 76.2 79.7 91.0 75.8 80.1 81.6 89.8 0.0 0.0 75.0 83.0 83.8 89.5 68.5 74.4 Token Distillation + NTP 77.3 82.6 68.4 73.8 75.8 86.3 72.3 79.9 77.3 86.1 62.3 69.1 65.2 75.8 65.4 74.6 70.5 78.5 Token Distillation + NTP 76.8 82.0 71.7 75.2 80.9 91.6 77.5 81.2 81.1 89.5 72.7 79.5 71.9 80.5 81.2 86.9 76.7 83.3
@@
Mistral-7B Llama3-8B Llama3-8B-i Qwen3-8B Avg. Original tokenization 69.5 69.4 72.1 81.7 73.2 Random (Hewitt, 2021) 51.0 46.1 51.5 62.1 52.7 Mean (Gee et al., 2022) 56.3 58.4 61.7 69.6 61.5 NTP Lampinen et al., 2018 64.7 67.0 70.1 81.2 70.8 ZeTT (Minixhofer et al., 2024) 66.1 61.3† 65.2† – – Token Distillation 68.5 68.9 72.9 81.5 72.9
@@ 5.1 Main Experiments
Note that even with its very competitive performance, Token Distillation does not fully attain the level of the original tokenization, which benefits from the massive pretraining of the original model. This is expected in our zero-shot tokenizer transfer setting without any further training (see Minixhofer et al., 2024), although remarkably, Token Distillation is able to outperform the original tokenization when adding French tokens to Llama3-8B-i without any training of the Transformer layers. We seek to directly investigate the embedding initialization quality and the advantages of our distillation-based objective compared to next-token prediction rather than the effects of further training. However, results can naturally be improved via further training; we argue that Token Distillation will provide the best starting point for such further processing. In Section 5.2, we additionally experiment with giving Token Distillation additional computational budget as well as a moderate continued training phase, which does enable Token Distillation to closely match the original tokenization’s performance while offering significant speedups due to reduced token counts. In the following, we analyze various aspects of the results in further detail.
Note that in Table 2, only for Llama3.2-3B, Token Distillation obtains results that are on par with random initialization. Llama3.2-3B(-i) are the only models in our lineup with tied embedding weights. Our objective does not explicitly enforce a bound on the norm of the new embedding, which in this case led to the failure mode of always generating a specific new embedding with very large norm. Note that this does not always happen for checkpoints with tied weights, as evidenced by the results for Llama3.2-3B-i. However, we additionally propose a simple modification to alleviate the occurrence of this issue by combining Token Distillation with the next-token prediction objective, which (implicitly) acts as a regularizer on the output embedding norm. In support of our argument that a subtoken attention distillation objective is superior to next-token prediction for learning new token embeddings, the combination via a sum of the two objectives (Token Distillation + NTP) generally yields worse results than Token Distillation alone. However, we can add a dynamic downweighting factor (see Section 4.2) to the next-token prediction objective (Token Distillation + NTP), which mostly alleviates the negative interference while keeping the regularizing effect.
@@
Vanilla NTP-tuning of all embeddings (denoted as tune all embeddings) yields disappointing results, even underperforming the subtoken mean initialization. We investigate this and observe a degradation of the original token embeddings. Note that for Token Distillation, we only optimize new token embeddings. Therefore, we also compare against the NTP baseline, where we similarly optimize only the new token embeddings. NTP outperforms NTP (tune all embeddings), supporting our analysis. However, even with this improvement, Token Distillation in turn still further surpasses NTP.
@@
ZeTT, which uses a pretrained hyper-network, is the strongest baseline we compare against. We only compare against ZeTT when a pretrained hyper-network by Minixhofer et al. (2024) is available444Minixhofer et al. (2024) do not provide a checkpoint trained specifically for the “-Instruct” variant of Llama3-8B but show that a transfer from the base version is possible.. We first note that ZeTT yields quite impressive results, outperforming even our optimized NTP baseline. Evidently, the pretraining of ZeTT’s hyper-network has internalized a better embedding prediction than the iterative gradient descent optimization on selected samples we employ for NTP. Nevertheless, our proposed method Token Distillation outperforms even ZeTT – without needing any hyper-network pretraining. It is important to note that at inference time, hyper-network based methods such as ZeTT are actually faster than Token Distillation (and NTP-based baselines), since only a single forward pass through the hyper-network per token is required. However, they require expensive pretraining for each new model and – in some cases – target domain. For example, ZeTT with a hyper-network pretrained only on English naturally underperforms on language adaptation to other languages – as this is out of distribution from the hyper-network training, whereas the multilingually pretrained variant available for one of the models performs more favorably (see Table 3). Also, we can actually use ZeTT-generated embeddings as a starting point and further tune them using our method for even better results. We analyze this possibility in Section 5.3.
5.2 Compression-Oriented Tokens and Continued Training
In our main experiments, we select new tokens based on frequently occurring whole words (yielding moderate sequence compression of 7–10%). In practice, however, vocabulary adaptation is often pursued for computational efficiency, which calls for selecting tokens that maximize sequence compression – even if these tokens are not whole words and are therefore inherently more challenging to initialize well (see our discussion in Appendix A). We therefore include two experiments that explicitly target this regime.
Compression-oriented tokens on French (Table 5).
We repeat French vocabulary adaptation, but (i) increase the number of snippets per new token from 25 to 100 (4), and (ii) select 5,000 new tokens using AdaptiVocab (Nakash et al., 2025). Importantly, AdaptiVocab does not restrict tokens to whole words, which increases sequence compression but also makes initialization harder (tokens are more ambiguous and more context-dependent than whole words). As a result, init-only scores in this experiment are not directly comparable to our main FrenchBench results in Table 3, which use a different (whole-word) token set. AdaptiVocab yields substantially stronger compression, reducing average token counts by about 20% on the evaluated FrenchBench datasets. Despite the increased difficulty of the token set, Token Distillation closely matches the original tokenization while providing the compression benefit.
Method Llama-8B-i ( tokens) Qwen3-8B ( tokens) Original tokenization 72.1 (0%) 81.7 (0%) Subtoken Mean 60.9 (-21%) 69.2 (-20%) NTP 66.5 (-21%) 79.1 (-20%) Token Distillation 72.2 (-21%) 81.5 (-20%)
Method Init only + Continued training tokens Original tokenization 47.4 46.9 0% Subtoken Mean 36.2 45.4 -36% NTP 40.6 46.1 -36% Token Distillation 44.2 47.5 -36%
Compression-oriented tokens + continued training on Arabic (Table 5).
We next consider a setting where added tokens come from a language that is very poorly supported in the source tokenizer. We add the top 1,000 Arabic tokens selected via AdaptiVocab to Mistral-7B, reducing token counts on Arabic benchmarks by 36%. Note that the reduction of token counts by 36% corresponds to compute savings of up to 50%, due to the quadratic scaling of attention with respect to sequence lengths (see Section B.4). Using the same increased initialization budget as in the previous experiment on French compression-oriented tokens (100 snippets per new token), Token Distillation already outperforms other initialization baselines (“Init only”) but does not yet match the original tokenization. We then run a moderate continued-training phase for vocabulary adaptation based on the methodology proposed by Nakash et al. (2025), which trains the embedding matrices as well as the first and last Transformer layers. After this continued training, Token Distillation matches or slightly exceeds the original tokenization while retaining the substantial token-count reduction. The original tokenization is not able to benefit from this moderate continued training, likely due to heavy over-fragmentation.